OpenStudy (anonymous):

x^2+9x+12, how would you know very quickly that this polynomial is prime ?

OpenStudy (anonymous):

no two numbers both multiply to 12 and add to 9

OpenStudy (anonymous):

really ?

OpenStudy (anonymous):

Well the factors of 12 are 1,12, 2,6, and 3,4. 1+12 = 13 2+6=8 3+4=7